# FastDraft: How to Train Your Draft

### v1

@@ 5.3 Continued pre-training for code
We investigate the continued pre-training (CP) method introduced by [Aryabumi et al., 2024] to refine our draft model on code-related tasks. In this approach, we extend the training of our pre-trained draft model using three distinct combinations of code and natural language datasets. Each combination incorporates 5BT of code data, along with varying amounts of natural language data: specifically, ${0,1,2.5}$BT. We then evaluate the resulting models on our benchmarks, including the HumanEval benchmark, and present the block efficiency results, $\tau^{\gamma}$, in Figure 2. Our findings indicate that mixed CP datasets lead to significantly higher block efficiency in natural language tasks compared to using CP with code alone. Additionally, block efficiency improves across natural language benchmarks, surpassing the performance of the base draft model that was trained solely on natural language, with only a slight decline observed in the Dolly benchmark. Notably, the base draft shows a substantial improvement on the HumanEval benchmark. The enhancement in the draft’s performance on natural language tasks with CP was anticipated, based on the findings of [Aryabumi et al., 2024]. However, unlike the conclusions presented in that study, which advocate for a balanced dataset of natural language and code, our results suggest that CP is the preferred approach for integrating code and natural language datasets when training a draft model.
@@
Table 2: AR results for fine-tuning with original data vs synthetic data
Sampling method Data source CNN-DailyMail TinyStories Dolly
$\gamma=3$ $\gamma=5$ $\gamma=3$ $\gamma=5$ $\gamma=3$ $\gamma=5$
Greedy None 0.350 0.246 0.295 0.196 0.259 0.174
Original 0.357 0.251 0.282 0.189 0.328 0.228
Target 0.378 0.270 0.296 0.201 0.370 0.262
Multinomial sampling None 0.311 0.221 0.227 0.184 0.245 0.163
Original 0.328 0.227 0.268 0.181 0.311 0.215
Target 0.339 0.242 0.286 0.193 0.352 0.234
5.4 Fine tuning with synthetic data vs original data
Since our objective is to achieve the best draft alignment with the target model rather than the best answers, we investigate the impact of using the target’s responses instead of the original dataset responses, as described in Section 4.2. We measure the draft’s performance across several tasks. In this experiment, we utilize the OIG-small-chip dataset for fine-tuning. We generate the target’s responses with multinomial sampling at temperature of 0.6. Token-level KD was not applied in this experiment. The results are presented in Table 2. An analysis of the data reveals that using the target’s answers leads to better draft alignment compared to using the original dataset answers across all tasks
Table 3: Knowledge distillation with KL-Divergence and TVD
Sampling method Loss CNN-DailyMail TinyStories Dolly
$\gamma=3$ $\gamma=5$ $\gamma=3$ $\gamma=5$ $\gamma=3$ $\gamma=5$
Greedy $\mathcal{L}_{CE}$ 0.378 0.270 0.296 0.201 0.370 0.262
$0.5\mathcal{L}_{CE}+0.5\mathcal{L}_{KL}$ 0.376 0.269 0.295 0.199 0.368 0.261
$0.5\mathcal{L}_{CE}+0.5\mathcal{L}_{TVD}$ 0.377 0.270 0.297 0.202 0.370 0.263
$\mathcal{L}_{KL}$ 0.374 0.266 0.294 0.198 0.371 0.261
$\mathcal{L}_{TVD}$ 0.384 0.274 0.301 0.204 0.362 0.254
Multinomial sampling $\mathcal{L}_{CE}$ 0.339 0.242 0.286 0.193 0.352 0.234
$0.5\mathcal{L}_{CE}+0.5\mathcal{L}_{KL}$ 0.350 0.248 0.289 0.190 0.346 0.235
$0.5\mathcal{L}_{CE}+0.5\mathcal{L}_{TVD}$ 0.356 0.243 0.297 0.194 0.347 0.244
$\mathcal{L}_{KL}$ 0.355 0.235 0.280 0.188 0.343 0.237
$\mathcal{L}_{TVD}$ 0.347 0.250 0.297 0.197 0.344 0.237
@@ 5.5 Fine tuning with knowledge distillation
To further enhance the alignment of our base draft model, we incorporate knowledge distillation alongside training on target-generated data, as discussed in Section 4.2. We experimented with various combinations of Cross Entropy (CE) loss, denoted as $\mathcal{L}_{CE}$, applied to the ground truth labels, and knowledge distillation losses, using the sparse logits collected from the target model: $\mathcal{L}_{KL}$ and $\mathcal{L}_{TVD}$. We utilize the same dataset we generated in Section 5.4 for this experiment. The results of these experiments are presented in Table 3. Our findings indicate that knowledge distillation in this setup does not provide a significant advantage over CE loss on the target-generated data. Although Total Variation Distance (TVD) has been shown to be negatively correlated with acceptance rates, in our results, it offers only slight benefit on some benchmarks over CE and Kullback–Leibler Divergence (KLD).
6 Results & reproducibility
Combining the findings from the ablation studies presented in section 5, we outline our comprehensive pipeline for draft pre-training and fine-tuning to optimize performance on key metrics for speculative decoding and demonstrate it by producing a draft for Phi-3-mini. To further illustrate reproducibility, we employ the same pipeline to produce a draft for Llama-3.1-8B-Instruct Dubey et al. [2024]. We exhibit a 50M draft for the Phi-3-mini and a 150M draft for Llama-3.1-8B. Drafts’ architecture details are presented in Appendix A.1. Considering both performance and resource efficiency, our findings from sections 5.2, 5.3 suggest pre-training over 5BT of FineWeb text data and continued pre-training on a mixture of 5BT code data from The Stack v2 and 2.5BT FineWeb text data as a best practice. For fine-tuning, we conclude from sections 5.4, 5.5 that sequence-level KD yields significant improvements while token-level KD benefits are not definitive, therefore, for FastDraft we only utilize sequence-level KD. We construct an alignment dataset for both Phi-3 and Llama-3.1 drafts by combining a number of synthetic datasets we generated with the appropriate target model as described in Section 5.1. Table 4 presents the AR improvements achieved through each stage of FastDraft generated with multinomial sampling. Results for greedy sampling are presented in Table 14 in the appendix. Using both sampling methods, we observe for datasets CNN-DailyMail, TinyStories, and Dolly, a substantial AR increase following fine tuning, with instruction-following dataset Dolly exhibiting a ~10% AR increase. For HumanEval the primary performance gains stem from pre-training on code domain knowledge during continued pre-training. These models achieve a MBSU of ~2x for natural language tasks and ~3x for code completion tasks.
The high quality of models like Phi-3-mini and Llama-3.1-8B, combined with advancements in edge device hardware, e.g. laptops and desktops, has made local deployment of LLMs increasingly appealing, though still challenging. Typical use cases for locally deployed LLMs share similar characteristics, they are designed for single-user interactions where only one query is processed at a time. This setup presents an excellent opportunity to apply speculative decoding to accelerate LLM generation with minimal resource demands. To evaluate this claim, we benchmark Phi-3-mini and Llama-3.1-8B using our FastDraft drafts on the latest Intel® Core Ultra series 200 (Lunar Lake) processors, with tests conducted on the CNN-DailyMail and HumanEval datasets. Both target and draft models are deployed on the chip’s integrated GPU using Intel® OpenVINO. We observe an average TPOT speedup of 1.5x for summarization tasks and 2x for code completion. Additionally, memory transaction efficiency improves, with up to 3x reduction in memory bandwidth usage, as indicated by the MBSU metric in these trials. Memory transactions can be highly energy-intensive [Horowitz, 2014], making the use of FastDraft drafts a promising approach for achieving substantial power savings on edge devices. These performance gains stem from the compact nature of our drafts, enabling them to operate in tandem with large target LLMs, and from their high acceptance rate, which reduces the number of calls to the target model. We believe FastDraft has the potential to unlock large language models inference on AI-PC and other edge-devices.
Model Type CNN-DailyMail TinyStories Dolly HumanEval
@@ 7 Conclusion & future work
In this paper, we introduced FastDraft, a novel approach for training and evaluating draft models for speculative decoding. Our results demonstrate that FastDraft facilitates the rapid training of high-quality, efficient draft models that are well-aligned with target models. We conducted a comprehensive ablation study, addressing various aspects of draft training, including pre-training data composition and draft-target alignment. Using our method, we successfully trained a highly efficient 50M-parameter draft model for Phi-3-mini, achieving up to a 67% acceptance rate and up to a 3x memory-bound speedup. We successfully demonstrated the scalability of FastDraft by training a draft for Llama-3.1-8B-Instruct, showing its effectiveness for this model as well. We validated our theoretical findings by benchmarking our drafts on the latest AI-PC hardware achieving a wall-clock time speedup of up to 2x. We hope that the findings of this work will inspire further research into efficient draft training, particularly focusing on the development of resource-efficient draft architectures and hardware-aware designs.
@@ A.4 Pre-training model size ablation
Draft size Data size Perplexity CNN-DailyMail TinyStories Dolly
@@ A.5 Pre-training on code ablation
Consistent with current best practices for pre-training large language models Dubey et al. [2024], Abdin et al. [2024], we experiment with pre-training the draft model on a mixture of code and natural language datasets with varying ratios of the domains. Similar to Aryabumi et al. [2024], we evaluate a balanced mixture of 5 billion token code and 5 billion token text datasets. Since our datasets are relatively small to begin with, emphasizing on text data can be beneficial for general language understanding across tasks. Accordingly, we also experiment with mixed datasets featuring 4 billion tokens of text and 6 billion tokens of code, as well as 7 billion tokens of text and 3 billion tokens of code.
@@
Continued pre-training variants with CP mixed dataset of code and text were able to obtain 1-4% higher acceptance rate on CNNT-DailyMail, Tinystories and Dolly datasets, over greedy and multinomial sampling methods (see Table 9, 11) compared to CP datasets of only code or only text. These models have been trained on 1-2.5 billion more tokens compared to pre-trained models on mixed data, but as section 5.2 suggests, the performance gain likely stems from the choice of pre-training strategy rather than additional data. Surprisingly, mixed datasets with higher proportion of text yield better results on HumanEval, a code evaluation dataset, with little to no improvement on natural language tasks 12, 13. This strengths our findings in section 5.2 that additional text data doesn’t contribute to our draft performance beyond a certain point.
@@
CP Base dataset CP dataset CP dataset CNN-DailyMail Tinystories Dolly
@@
CP Base dataset CP mix dataset CP mix dataset CNN-DailyMail Tinystories Dolly
@@
Mix dataset Mix dataset CNN-DailyMail Tinystories Dolly HumanEval
@@
Mix data Mix data CNN-DailyMail Tinystories Dolly HumanEval
@@
Appendix B FastDraft additional results
Table 4 summerizes results of each stage of the FastDraft scheme using greedy sampling.
Model Type CNN-DailyMail TinyStories Dolly HumanEval

### v2

@@ 5.3 Continued pre-training for code
We investigate the continued pre-training (CP) method introduced by (Aryabumi et al., 2024) to refine our draft model on code-related tasks. In this approach, we extend the training of our pre-trained draft model using three distinct combinations of code and natural language datasets. Each combination incorporates 5BT of code data, along with varying amounts of natural language data: specifically, ${0,1,2.5}$BT. We then evaluate the resulting models on our benchmarks, including the HumanEval benchmark, and present the block efficiency results, $\tau^{\gamma}$, in Figure 2. Our findings indicate that mixed CP datasets lead to significantly higher block efficiency in natural language tasks compared to using CP with code alone. Additionally, block efficiency improves across natural language benchmarks, surpassing the performance of the base draft model that was trained solely on natural language, with only a slight decline observed in the Dolly benchmark. Notably, the base draft shows a substantial improvement on the HumanEval benchmark. The enhancement in the draft’s performance on natural language tasks with CP was anticipated, based on the findings of (Aryabumi et al., 2024). However, unlike the conclusions presented in that study, which advocate for a balanced dataset of natural language and code, our results suggest that CP is the preferred approach for integrating code and natural language datasets when training a draft model.
@@
5.4 Fine tuning with synthetic data
Since our objective is to achieve the best draft alignment with the target model rather than the quality of generation, we investigate the impact of using the target’s responses instead of the original dataset responses, as described in Section 4.2. We measure the draft’s performance across several tasks. In this experiment, we utilize the OIG-small-chip dataset for fine-tuning. We generate the target’s responses with multinomial sampling at temperature of 0.6. Token-level KD was not applied in this experiment. The results are presented in Table 2. An analysis of the data reveals that using the target’s answers leads to better draft alignment compared to using the original dataset answers across all tasks.
Sampling Data source CNN-DM TS Dolly
Greedy None 0.350 0.295 0.259
Original 0.357 0.282 0.328
Target 0.378 0.296 0.370
Multinomial None 0.311 0.227 0.245
Original 0.328 0.268 0.311
Target 0.339 0.286 0.352
Table 2: AR results for fine-tuning with original data vs synthetic data. Results for block size $\gamma=3$
@@ 5.5 Fine tuning with knowledge distillation
To further enhance the alignment of our base draft model, we incorporate KD alongside training on target-generated data, as discussed in Section 4.2. We experimented with various combinations of Cross Entropy (CE) loss, denoted as $\mathcal{L}_{CE}$, applied to the ground truth labels, and KD losses, using the sparse logits collected from the target model: $\mathcal{L}_{KL}$ and $\mathcal{L}_{TVD}$. We utilize the same dataset we generated in Section 5.4 for this experiment. The results of these experiments are presented in Table 3. Our findings indicate that KD in this setup doesn’t provide significant advantage over CE loss on the target-generated data. Although Total Variation Distance (TVD) was shown to negatively correlate with acceptance rates, in our results, it offers only slight benefit on some benchmarks over CE and Kullback–Leibler Divergence (KLD).
Sampling Loss CNN-DM TS Dolly
Greedy $\mathcal{L}_{CE}$ 0.378 0.296 0.370
$\frac{1}{2}\mathcal{L}_{CE}+\frac{1}{2}\mathcal{L}_{KL}$ 0.376 0.295 0.368
$\frac{1}{2}\mathcal{L}_{CE}+\frac{1}{2}\mathcal{L}_{TVD}$ 0.377 0.297 0.370
$\mathcal{L}_{KL}$ 0.374 0.294 0.371
$\mathcal{L}_{TVD}$ 0.384 0.301 0.362
Multinomial $\mathcal{L}_{CE}$ 0.339 0.286 0.352
$\frac{1}{2}\mathcal{L}_{CE}+\frac{1}{2}\mathcal{L}_{KL}$ 0.350 0.289 0.346
$\frac{1}{2}\mathcal{L}_{CE}+\frac{1}{2}\mathcal{L}_{TVD}$ 0.356 0.297 0.347
$\mathcal{L}_{KL}$ 0.355 0.280 0.343
$\mathcal{L}_{TVD}$ 0.347 0.297 0.344
Table 3: KD with KL-Divergence and TVD. Results for block size $\gamma=3$
Model Type CNN-DM TS Dolly HumanEval
@@
5.6 Hardware-aware draft
The high quality of models like Phi-3-mini and Llama-3.1-8B, combined with advances in edge device hardware, has made local deployment of LLMs increasingly appealing, yet still challenging. Typical use cases for locally deployed LLMs share similar characteristics, they are designed for single-user interactions where only one query is processed at a time. This setup presents an excellent opportunity to apply speculative decoding to accelerate LLM generation with minimal resource demands.
The Phi3-mini 50M uses the original Transformers dimensions, however, due to the strict latency requirements of draft models, their architectural design warrants special attention, as their performance may vary depending on the hardware platform. We conduct a series of experiments to examine the performance of draft models derived from different architectural choices. Our focus is on altering the depth of the models by changing the number of layers or the width of the models by modifying the hidden dimension. The target system for our experiments is the latest Intel® Core Ultra series 200 (Lunar Lake) processors coupled with Intel® OpenVINO runtime engine.
First, we analyze the effect of scaling the width and depth of the draft on latency of the draft model on the target hardware. We fix either the hidden dimension or the model depth to match the corresponding values pf Phi3-mini 50M draft and modify the other parameter to observe how latency scales with wider or deeper models. As shown in Figure 3, our benchmarking results suggest that increasing the number of layers has the most significant impact on latency, while increasing the hidden dimension leads to only a slight increase.
Next, we evaluate the impact of the draft architecture under a fixed parameter budget4 4 4 We excluded embedding table size from the model parameter budget, as it does not strain memory bandwidth. by assessing the AR and estimated speedup achieved when varying the width and depth of the draft. By redistributing parameters between these dimensions, we aimed to identify configurations that optimize the trade-off between speculation quality and computational efficiency. We pre-trained multiple draft models for Phi-3-mini, varying the number of layers and hidden dimension sizes while maintaining a fixed parameter count. We then measured their AR on a target dataset and recorded their latency on our target hardware to estimate the speedup achieved through SD via Equation 2. Figure 4 presents the results for the CNN-DM dataset. Our findings indicate that increasing depth does not improve the AR, while excessively shallow drafts degrade performance. Moreover, the AR–latency trade-off follows an inverted U-shaped speedup curve with a broad base, consistent with our earlier experimental results. The results for the TS and Dolly datasets, presented in Appendix B, exhibited the same pattern.
Figure 3: Latency of models obtained by varying either a hidden dimension or a number of layers
Figure 4: Estimated speedup on CNN-DM for models with constant parameter budget. L, H denotes the number of layers and hidden dimension size
6 Results & reproducibility
Combining the findings from the ablation studies presented in Section 5, we outline our comprehensive pipeline for draft pre-training and fine-tuning to optimize performance on key metrics for speculative decoding and demonstrate it by producing a draft for Phi-3-mini. To further illustrate reproducibility, we employ the same pipeline to produce a draft for Llama-3.1-8B-Instruct Dubey et al. (2024). We present a 50M draft for the Phi-3-mini and a 150M draft for Llama-3.1-8B. Drafts’ architecture details are presented in Appendix A.1. Considering both performance and resource efficiency, our findings from sections 5.2, 5.3 suggest pre-training over 5BT of FineWeb text data and continued pre-training on a mixture of 5BT code data from The Stack v2 and 2.5BT FineWeb text data as a best practice. For fine-tuning, we conclude from sections 5.4, 5.5 that sequence-level KD yields significant improvements while token-level KD benefits are not definitive, therefore, for FastDraft we only utilize sequence-level KD. We construct an alignment dataset for both Phi-3 and Llama-3.1 drafts by combining a number of synthetic datasets we generated with the appropriate target model as described in Section 5.1. Table 4 presents the AR improvements achieved through each stage of FastDraft generated with multinomial sampling. Results for greedy sampling are presented in Table 17 in the Appendix. Using both sampling methods, we observe for datasets CNN-DM, TS, and Dolly, a substantial AR increase following fine tuning, with instruction-following dataset Dolly exhibiting an increase of ~10% in AR. For HumanEval the primary performance gains stem from pre-training on code domain knowledge during continued pre-training. These models achieve a MBSU of ~2x for natural language tasks and ~3x for code completion tasks.
Furthermore, we demonstrate the real-world wall-clock time speedup achieved with FastDraft on the Intel® Core Ultra series 200 as detailed in Section 5.6. We observe an average TPOT speedup of 1.5x for summarization tasks and 2x for code completion. Additionally, memory transaction efficiency improves, with up to 3x reduction in memory bandwidth usage, as indicated by the MBSU metric in these trials. Memory transactions can be highly energy-intensive (Horowitz, 2014), making the use of FastDraft drafts a promising approach for achieving substantial power savings on edge devices. These performance gains stem from the compact nature of our drafts, enabling them to operate in tandem with large target LLMs. We believe FastDraft has the potential to unlock LLMs inference on AI-PC and other edge-devices.
@@ 7 Conclusion & future work
In this paper, we introduced FastDraft, a novel approach for training and evaluating draft models for speculative decoding. Our results demonstrate that FastDraft facilitates the rapid training of high-quality, efficient draft models that are well-aligned with target models. We conducted a comprehensive ablation study to examine various aspects of draft training, including pre-training data composition and draft-target alignment. Using our method, we successfully trained a highly efficient 50M-parameter draft model for Phi-3-mini, achieving an acceptance rate of up to 67% and up to a 3x memory-bound speedup. We further demonstrated the scalability of FastDraft by training a draft for Llama-3.1-8B-Instruct, underscoring its effectiveness for larger models. We validated our theoretical findings by benchmarking our drafts on the latest AI-PC HW achieving a wall-clock time speedup of up to 2x. We hope our findings will inspire additional research on efficient draft training, particularly focusing on the development of resource-efficient draft architectures and hardware-aware designs.
8 Limitations
Our training recipe for draft models has been validated exclusively in English, and the reported acceptance rates and speedup benefits may not generalize to non-English languages due to their distinct syntactic structures and linguistic characteristics. Future work should investigate the effectiveness of our approach across different languages and develop language-specific modifications to optimize performance for non-English applications.
While prior work has demonstrated that speculative decoding with multiple candidate sequences can achieve higher acceptance rates and greater speedup compared to single-sequence approaches (Li et al., 2024; Cai et al., 2024), we restricted our implementation to single-sequence speculation due to computational constraints. Multiple-sequence speculation requires substantial parallel compute resources to generate and validate multiple candidate sequences simultaneously. Given our focus on AI-PC, we prioritized developing methods that could run effectively on consumer HW.
Our work primarily investigates optimal training strategies for draft models that share the same architectural design as their target models. While this focused scope allowed us to deeply explore the relationship between architecturally identical draft and target pairs, it also presents notable limitations. First, our findings may not generalize well to scenarios where draft and target models employ different architectures, as the dynamics of knowledge transfer and performance optimization could vary significantly in such cases. Additionally, while our proposed training recipe proved effective for architecturally identical pairs, it’s possible that combining our approach with more efficient draft architectures could yield even better results—potentially offering improved speed-quality trade-offs that we have not explored in this work.
@@ A.4 Pre-training model size ablation
Draft size Data size Perplexity CNN-DM TS Dolly
@@ A.5 Pre-training on code ablation
Consistent with current best practices for pre-training large language models Dubey et al. (2024); Abdin et al. (2024), we experiment with pre-training the draft model on a mixture of code and natural language datasets with varying ratios of the domains. Similar to Aryabumi et al. (2024), we evaluate a balanced mixture of 5 billion token code and 5 billion token text datasets. Since our datasets are relatively small to begin with, emphasizing on text data can be beneficial for general language understanding across tasks. Accordingly, we also experiment with mixed datasets featuring 4 billion tokens of text and 6 billion tokens of code, as well as 7 billion tokens of text and 3 billion tokens of code.
@@
Continued pre-training variants with CP mixed dataset of code and text were able to obtain 1-4% higher acceptance rate on CNN-DM, TS and Dolly datasets, over greedy and multinomial sampling methods (see Table 9, 11) compared to CP datasets of only code or only text. These models have been trained on 1-2.5 billion more tokens compared to pre-trained models on mixed data, but as section 5.2 suggests, the performance gain likely stems from the choice of pre-training strategy rather than additional data. Surprisingly, mixed datasets with higher proportion of text yield better results on HumanEval, a code evaluation dataset, with little to no improvement on natural language tasks 12, 13. This strengths our findings in section 5.2 that additional text data doesn’t contribute to our draft performance beyond a certain point.
@@
CP Base dataset CP dataset CP dataset CNN-DM TS Dolly
@@
CP Base dataset CP mix dataset CP mix dataset CNN-DM TS Dolly
@@
Mix dataset Mix dataset CNN-DM TS Dolly HumanEval
@@
Mix data Mix data CNN-DM TS Dolly HumanEval
@@
Appendix B Hardware-aware draft details and results
Tables 14 and 15 contain the full list of configuration used in HW-aware experiments. Table 16 lists the configurations used for estimating the speedup of architectures with varying depth or width satisfying a constant parameter budget.
Figures 5 and 6 present the results for the constant budget pre-trained drafts evaluated on the TS and Dolly datasets.
# Params (M) Hidden size Intermediate size # Layers #Attention heads # Key-value heads
27.91 512 1408 4 8 8
38.95 640 1792 4 10 10
37.10 512 1408 7 8 8
42.67 768 2048 4 12 12
43.23 512 1408 9 8 8
80.32 1024 2816 4 16 16
80.00 512 1408 21 8 8
114.78 1280 3456 4 20 20
113.70 512 1408 32 8 8
34.04 512 1408 6 8 8
48.64 640 1792 6 10 10
49.36 512 1408 11 8 8
64.00 768 2048 6 12 12
64.68 512 1408 16 8 8
104.83 1024 2816 6 16 16
104.51 512 1408 29 8 8
152.60 1280 3456 6 20 20
153.53 512 1408 45 8 8
Table 14: Phi-3-mini draft configurations for latency benchmarks. #Params values exclude the embeddings table parameters
Name Hidden size Intermediate size # Layers #Attention heads # Key-value heads
Phi3-draft-256 256 768 32 4 4
Phi3-draft-384 384 1024 12 6 6
Phi3-draft-448 448 1280 8 7 7
Phi3-draft-576 576 1536 4 9 9
Phi3-draft-640 640 1792 3 10 10
Phi3-draft-768 768 2048 2 12 12
Phi3-draft-base 512 1408 6 8 8
Table 15: Configurations of models sharing the parameter budget with the Phi3-mini 50M draft configuration
Model CNN-DM TS Dolly
AR Speedup AR Speedup AR Speedup
Phi3-mini-256 0.3319 1.2481 0.3182 1.2223 0.2681 1.1283
Phi3-mini-384 0.3170 1.5616 0.3071 1.5379 0.2097 1.3039
Phi3-mini-448 0.3255 1.6222 0.3009 1.5616 0.2177 1.3568
Phi3-mini-512 0.3277 1.6941 0.3021 1.6284 0.2351 1.4568
Phi3-mini-576 0.3230 1.7066 0.2905 1.6220 0.2222 1.4444
Phi3-mini-640 0.3104 1.6918 0.2836 1.6212 0.2251 1.4676
Phi3-mini-768 0.2236 1.4756 0.2713 1.6020 0.2006 1.4149
Table 16: AR and Estimated speedup results for models obtained by modifying depth and width to stay within the same parameter budget as Phi3-mini-base draft model
Figure 5: Estimated speedup on TS for models with
constant parameter budget. L, H denotes the number
of layers and hidden dimension size
Figure 6: Estimated speedup on Dolly for models with
constant parameter budget. L, H denotes the number
of layers and hidden dimension size
Figure 7: First and second token latencies of models obtained by varying either a hidden dimension or a number of layers
Appendix C FastDraft additional results
Table 4 summarizes results of each stage of the FastDraft scheme using greedy sampling.
Model Type CNN-DM TS Dolly HumanEval